Find three consecutive even integers such that if the largest integer is subtracted from four times the smallest, the result is 6 more than twice the middle integer. (Enter your answers as a comma-separated list.)
Answer by ewatrrr(24785)

three consecutive even integers
x
x+2
x+4
4x - (x+4) = 2(x+2) + 6
x = 14
The three consecutive even integers are: 14, 16, 18
checking..
56 - 18 = 38 = 32 + 6